Two baseball teams engage in a playoff series in which the series is won by the first team to win 4 games. If each team is equally likely to win any game, what is the expected number of games played?
Answer #1

P(game is played for 4 games)=P(X=4)=P(first team wins first 4 games)+P(2nd team wins first 4 games)

=2*(0.5)4 =0.125

P(X=5)=P(first team wins 3 out of first 4 games and wins 5th game)+P(2nd team wins 3 out of first 4 games and wins 5th game)

=2*4C3(0.5)4*(0.5)=0.25

P(X=6)=2*5C3(0.5)4*(0.5)2=0.3125

P(X=7)=2*6C3(0.5)4*(0.5)3=0.3125

hence Expected number of games played =E(X)=4*0.125+5*0.25+6*0.3125+7*0.3125=5.8125